# A First Introduction to Cooperative Multi-Agent Reinforcement Learning

### v4

@@ 1.1 Overview
Decentralized training and execution methods make the fewest assumptions and are often simple to implement. In fact, any single-agent RL method can be used for DTE by just letting each agent learn separately. Naturally, these approaches have their respective advantages and disadvantages, which will be discussed. It is worth noting that DTE is required if no centralized training phase is available (e.g., through a centralized simulator), requiring all agents to learn during online interactions without prior coordination. DTE methods can be applied in cooperative, competitive, or mixed cases. MARL methods can be further categorized into value-based and policy gradient methods. Value-based methods (e.g., Q-learning) learn a value function and then choose actions based on those values. Policy gradient methods learn an explicit policy representation and attempt to improve the policy in the direction of the gradient. Both classes of methods are widely used in MARL.
@@ 3.2.3 Deep extensions, issues, and fixes
Again, the code can be adjusted to the episodic/indefinite-horizon case by replacing the horizon with terminal states or the infinite-horizon case by removing the loop over episodes. The pseudocode, for this approach and throughout the text, is written to be simple and highlight the main algorithmic components. There are many implementation tricks, especially in the deep case, that are not included.
@@ 4.2.2 QMIX
then, [${Q}_{i}$] satisfy IGM for $Q$ at $\mathbf{h}{}$.
@@ 4.2.6 The use of state in factorization methods
Many of the current methods use inconsistent notation about the inclusion of state (except QMIX). As a result, the theory is often developed without considering the state input. It turns out that, unlike the policy gradient case discussed in Section 4.3.6, using the state doesn’t introduce bias and is thus theoretically sound (Marchesini et al., 2024). The intuition is that the state is additional information in these cases. It doesn’t replace the history, but potentially augments it (similar to a history-state critic in Section 4.3.6). Furthermore, actions are not chosen based on the state but only the local, history-dependent Q-values, ${Q}_{i}$. Nevertheless, in implementations of algorithms such as QPLEX and weighted QMIX, the weights only take the state as input and not the history. This replacement of history with state in these contexts limits the representational power of the weights in partially observable problems. As a result, using the state as input in QPLEX (instead of the joint history) prevents it from being able to represent the full IGM class of functions. It still may be beneficial to use the state instead of the joint history but exploring what additional information to use and in what way is an interesting open question.
@@ 4.3.4 COMA
The COMA algorithm can also be an extension of Algorithm 6. The critic update can be calculated in the same way (using the Q-values from the network described above) but the agent-specific advantage is incorporated in the actor update. That is, during the actor update (starting on line 16), the agent-advantage ${A}_{i}(\mathbf{h}{},\mathbf{a})$ is calculated for the given agent and ${A}_{i}$ is used in the actor gradient estimate instead of $\hat{\mathbf{Q}}$.
@@ 4.3.5 MAPPO
where $r_{{\psi}_{i},i}=\frac{\pi_{{\psi}_{i}}({a}_{i}|{h}_{i})}{\pi_{\psi_{i,old}}({a}_{i}|{h}_{i})}$ [^10] Here, I consider the stochastic (not continuous) action case and for simplicity do not include GAE, mini-batching, or a policy entropy term (which is common in policy gradient methods to improve exploration). Also, the paper uses a local advantage loss, ${A}_{i}$, but this value would not be different than $\mathbf{A}$ in a general Dec-POMDP if it is updated using $\mathbf{A}$ and $r$ since these quantities would be the same for all agents. Finally, policies over single observations rather than histories are used, which as noted before is typically not sufficient for partially observable problems (general Dec-POMDPs) and state-based critics are used, which is incorrect as discussed below. Maximizing $\frac{\pi_{{\psi}_{i}}({a}_{i}|{h}_{i})}{\pi_{\psi_{i,old}}({a}_{i}|{h}_{i})}\mathbf{A}$ seeks to maximally improve the new policy, $\pi_{{\psi}_{i}}$, compared to the old policy, $\pi_{\psi_{i,old}}$, by reweighing the advantage using an importance sampling ratio (considering the advantage was calculated using the old policy).
@@
with $\hat{A}_{i}=r_{t}+\gamma\hat{V}_{i}({h}_{i,t+1})-\hat{V}_{i}({h}_{i,t})$. The IPPO critic loss becomes:
@@ 4.3.9 Other centralized critic methods
Many other methods use centralized critics. For instance, Iqbal and Sha (2019) use attention for determining which other agent information to include in the centralized critic. Qatten (Yang et al., 2020) extends QMIX (and weighted QMIX) to add structure to the mixing network, where the structure is inspired by a linear approximation of the Q-function and one set of weights is leaned using attention. A wide range of other approaches have been developed but are not included in this introduction to the area.
@@ 4.4.1 Adding centralized information to decentralized methods
A number of methods allow agents to take turns learning. For instance, DTE methods can be used but all agents are fixed (not learning) except for one. The learning agent learns until convergence, generating a best-response to the other agent policies. If this process continues until no agents can further improve their polices, the result is a Nash equilibrium (Banerjee et al., 2012; Su et al., 2024). In limited settings with additional strong assumptions (e.g., deterministic environments, full observability, additional coordination mechanisms to ensure coordinated policies), such methods can potentially converge to an optimal solutions (Lauer and Riedmiller, 2000; Jiang and Lu, 2023, 2022). While these methods are sometimes called ‘decentralized’ since coordination and communication is need during learning that is not available during execution, I consider them to be CTDE.
@@ Acknowledgements
I thank Andrea Baisero, Shuo Liu, Roi Yehoshua, and the other members of my Lab for Learning and Planning in Robotics (LLRP) for reading my (very) rough drafts and providing comments that helped improve the document as well as Frans Oliehoek for the many discussions about these topics and the resulting text.

### v5

# An Initial Introduction to Cooperative Multi-Agent Reinforcement Learning

@@ 1.1 Overview
Decentralized training and execution methods make the fewest assumptions and are often simple to implement. In fact, any single-agent RL method can be used for DTE by just letting each agent learn separately. Naturally, these approaches have their respective advantages and disadvantages, which will be discussed. It is worth noting that DTE is required if no centralized training phase is available (e.g., through a centralized simulator), requiring all agents to learn during online interactions without prior coordination. DTE methods can be applied in cooperative, competitive, or mixed cases.
MARL methods can be further categorized into value-based and policy gradient methods. Value-based methods (e.g., Q-learning) learn a value function and then choose actions based on those values. Policy gradient methods learn an explicit policy representation and attempt to improve the policy in the direction of the gradient. Both classes of methods are widely used in MARL.
@@ 1.3.1 Value-based methods
Pseudocode for a version of DRQN is shown in Algorithm 1. The algorithm uses a learning rate, $\alpha$, some form of exploration, which in this case is $\epsilon$-greedy, and $Q$ is estimated using an RNN parameterized by $\theta$ along with using a target network, $\theta^{-}$, and a replay buffer, $\mathcal{D}$. Episodes (either a fixed set or until convergence) are generated by interacting with the environment and these are added to the replay buffer. This buffer typically has a fixed size and new data replaces old data when the buffer is full. Episodes are then be sampled from the replay buffer (a single episode is sampled here but it can also be a minibatch of episodes). In order to have the correct internal state, RNNs are trained sequentially. Training is shown here from the beginning of the episode until the end, calculating the internal state and the loss at each step and updating $\theta$ using gradient descent. The target network $\theta^{-}$ is updated periodically (every $C$ episodes in the code) by copying the parameters from $\theta$.
The code is written from the finite-horizon perspective but it can be adjusted to the episodic or indefinite-horizon case by replacing the horizon with terminal states or the infinite-horizon case by removing the loop over episodes. The pseudocode, for this approach and throughout the text, is written to be simple and highlight the main algorithmic components. There are many implementation tricks that are not included.
2:  Initialize network parameters $\theta$ and $\theta^{-}$ for $Q_{\theta}$ and $Q_{\theta}^{-}$
3:  $\mathcal{D}\leftarrow\emptyset$
4:  $e\leftarrow 1$episode index}
5:  for all episodes do
6:     $h\leftarrow\emptyset$initial history is empty}
7:     for $t=1$ to $\mathcal{H}$ do
8:        Choose $a$ at $h$ from $Q^{\theta}(h,\cdot)$ with exploration (e.g., $\epsilon$-greedy)
9:        See reward $r$, observation $\o$
10:        append $a,o,r$ to $\mathcal{D}^{e}$
11:        $h\leftarrow hao$update RNN state of the network}
12:     end for
13:     sample an episode from $\mathcal{D}$
14:     for $t=1$ to $\mathcal{H}$ do
15:        $h\leftarrow\emptyset$
16:        $a,o,r\leftarrow\mathcal{D}^{e}(t)$
17:        $h^{\prime}\leftarrow hao$
18:        $y=r+\gamma\max_{a^{\prime}}Q^{\theta^{-}}(h^{\prime},a^{\prime})$
19:        Perform gradient descent on parameters $\theta$ with learning rate $\alpha$ and loss: $\big{(}y-Q^{\theta}(h,a)\big{)}^{2}$
20:        $h\leftarrow h^{\prime}$
21:     end for
22:     if $e\mod C=0$ then
23:        $\theta^{-}\leftarrow\theta$
24:     end if
25:     $e\leftarrow e+1$
26:  end for
27:  return  $Q$
Algorithm 1 DRQN (finite-horizon*)
@@ 1.3.2 Policy gradient methods
Pseudocode for a simple version of partially observable (advantage) actor-critic is shown in Algorithm 2. This approach is often referred to as advantage actor-critic since it uses the advantage $A(h,a)=Q(h,a)-V(h)$, representing the difference between the Q-value for a given action and the V-value at that history. In the algorithm, $A(h,a)$ isn’t explicitly stored since the TD error is used as an approximation for the advantage—$\hat{Q}(h,a)$ is approximated with $r_{t}+\gamma\hat{V}(h_{t+1})$, which is true in expectation. $\delta_{t}$ can be thought of as a sample of $\hat{Q}(h_{t},a_{t})-\hat{V}(h_{t})=\hat{A}(h_{t},a_{t})$. Otherwise, actor and critic models are initialized and then data is generated from episodes. On-policy updates are for the value function and the actor and critic are updated as discussed above.
Algorithm 2 Advantage Actor-Critic (A2C) (finite-horizon)
0:  Actor model $\pi(a|h)$, parameterized by $\psi$
0:  Critic model $\hat{V}(h)$, parameterized by $\theta$
1:  Initialized $\alpha$, and $\beta$ learning rates for actor and critic)
2:  for all episodes do
3:     $h_{0}\leftarrow\emptyset$Empty initial history}
4:     for $t=0$ to $\mathcal{H}-1$ do
5:        Choose $a_{t}$ at $h_{t}$ from $\pi(a|h_{t})$
6:        See reward ${r}_{t}$ and observation $o_{t}$
7:        $h_{t+1}\leftarrow h_{t}a_{t}o_{t}${Append new action and obs to previous history}
8:        Compute value TD error: $\delta_{t}\leftarrow{r}_{t}+\gamma\hat{V}(h_{t+1})-\hat{V}(h_{t})$
9:        Compute actor gradient estimate: $\gamma^{t}\delta_{t}\nabla\log\pi(a_{t}|h_{t})$
10:        Update actor parameters $\psi$ using gradient estimate (e.g., $\psi\leftarrow\psi+\alpha\gamma^{t}\delta_{t}\nabla\log\pi(a_{t}|h_{t})$)
11:        Compute critic gradient estimate: $\delta_{t}\nabla\hat{V}_{i}(h_{t})$
12:        Update critic parameters $\theta$ using gradient estimate (e.g., $\theta\leftarrow\theta+\beta\gamma\delta_{t}\nabla\hat{V}(h_{t}))$
13:     end for
14:  end for
@@ 3.2.3 Deep extensions, issues, and fixes
Again, the code can be adjusted to the episodic/indefinite-horizon case by replacing the horizon with terminal states or the infinite-horizon case by removing the loop over episodes.
@@ 4.2.2 QMIX
then, [${Q}_{i}$] satisfy IGM for $\mathbf{Q}$ at $\mathbf{h}{}$.
@@ 4.2.6 The use of state in factorization methods
Many of the current methods use inconsistent notation about the inclusion of state (except QMIX). As a result, the theory is often developed without considering the state input. It turns out that, unlike the policy gradient case discussed in Section 4.3.6, using the state doesn’t introduce bias and is thus theoretically sound (Marchesini et al., 2025). The intuition is that the state is additional information in these cases. It doesn’t replace the history, but potentially augments it (similar to a history-state critic in Section 4.3.6). Furthermore, actions are not chosen based on the state but only the local, history-dependent Q-values, ${Q}_{i}$. Nevertheless, in implementations of algorithms such as QPLEX and weighted QMIX, the weights only take the state as input and not the history. This replacement of history with state in these contexts limits the representational power of the weights in partially observable problems. As a result, using the state as input in QPLEX (instead of the joint history) prevents it from being able to represent the full IGM class of functions. It still may be beneficial to use the state instead of the joint history but exploring what additional information to use and in what way is an interesting open question.
4.2.7 Other value function factorization methods
Qatten (Yang et al., 2020) extends QMIX (and weighted QMIX) to add structure to the mixing network, where the structure is inspired by a linear approximation of the Q-function and one set of weights is leaned using attention. A wide range of other approaches have been developed but are not included in this introduction to the area.
@@ 4.3.4 COMA
The COMA algorithm can also be an extension of Algorithm 8. The critic update can be calculated in the same way (using the Q-values from the network described above) but the agent-specific advantage is incorporated in the actor update. That is, during the actor update (starting on line 16), the agent-advantage ${A}_{i}(\mathbf{h}{},\mathbf{a})$ is calculated for the given agent and ${A}_{i}$ is used in the actor gradient estimate instead of $\hat{\mathbf{Q}}$: $\gamma^{t}{A}_{i}(\mathbf{h}{}_{t},\mathbf{a}_{t})\nabla_{\!{\psi}_{i}}\log{\pi}_{i}(a_{i,t}|h_{i,t})$.
@@ 4.3.5 MAPPO
where $r_{{\psi}_{i},i}=\frac{\pi_{{\psi}_{i}}({a}_{i}|{h}_{i})}{\pi_{\psi_{i,old}}({a}_{i}|{h}_{i})}$. [^10] Here, I consider the stochastic (not continuous) action case and for simplicity do not include GAE, mini-batching, or a policy entropy term (which is common in policy gradient methods to improve exploration). Also, the paper uses a local advantage loss, ${A}_{i}$, but this value would not be different than $\mathbf{A}$ in a general Dec-POMDP if it is updated using $\mathbf{A}$ and $r$ since these quantities would be the same for all agents. Finally, policies over single observations rather than histories are used, which as noted before is typically not sufficient for partially observable problems (general Dec-POMDPs) and state-based critics are used, which is incorrect as discussed below. Maximizing $\frac{\pi_{{\psi}_{i}}({a}_{i}|{h}_{i})}{\pi_{\psi_{i,old}}({a}_{i}|{h}_{i})}\mathbf{A}$ seeks to maximally improve the new policy, $\pi_{{\psi}_{i}}$, compared to the old policy, $\pi_{\psi_{i,old}}$, by reweighing the advantage using an importance sampling ratio (considering the advantage was calculated using the old policy).
@@
with $r_{{\psi}_{i},i}=\frac{\pi_{{\psi}_{i}}({a}_{i}|{h}_{i})}{\pi_{\psi_{i,old}}({a}_{i}|{h}_{i})}$ as before and $\hat{A}_{i}=r_{t}+\gamma\hat{V}_{i}({h}_{i,t+1})-\hat{V}_{i}({h}_{i,t})$. The IPPO critic loss becomes:
@@ 4.3.9 Other centralized critic methods
Many other methods use centralized critics. For instance, Iqbal and Sha (2019) use attention for determining which other agent information to include in the centralized critic. Other approaches are related to the ones presented here but some form of MAPPO is still the most widely used centralized critic method.
@@ 4.4.1 Adding centralized information to decentralized methods
A number of methods allow agents to take turns learning. For instance, DTE methods can be used but all agents are fixed (not learning) except for one. The learning agent learns until convergence, generating a best-response to the other agent policies. If this process continues until no agents can further improve their polices, the result is a Nash equilibrium (Banerjee et al., 2012; Su et al., 2024). In limited settings with additional strong assumptions (e.g., deterministic environments, full observability, additional coordination mechanisms to ensure coordinated policies), such methods can potentially converge to optimal solutions (Lauer and Riedmiller, 2000; Jiang and Lu, 2023, 2022). While these methods are sometimes called ‘decentralized’ since coordination and communication is need during learning that is not available during execution, I consider them to be CTDE.
@@ Acknowledgements
I thank Andrea Baisero, Shuo Liu, Roi Yehoshua, and the other members of my Lab for Learning and Planning in Robotics (LLRP) for reading my (very) rough drafts and providing comments that helped improve the document as well as Frans Oliehoek for the many discussions about these topics and the resulting text. I also thank NSF CAREER Award 2044993 for funding this effort.